OpenStudy (anonymous):

Determine whether each statement is true or false. Explain. The number of permutations of two items from a data set is always two times the number of combinations when taking two objects at a time from the same data set.

OpenStudy (anonymous):

nP2 = 2(nC2) ?

OpenStudy (anonymous):

\[nP2 = \frac{n!}{(n-2)!}\] \[nC2 = \frac{n!}{(n-2)!*2!}\]

OpenStudy (anonymous):

it was False

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I'll let you figure it out. Check some test cases if you need to.